# thanks: These authors contributed equally to this workthanks: These authors contributed equally to this workthanks: These authors contributed equally to this work

### v1

@@ Abstract
From the above analysis, accurate EPC predictions can be expected within the KS-DFT framework if the SIE in the exchange-correlation approximation is minimized. This can be achieved by advancing up Jacob’s ladder of DFT Perdew and Schmidt (2001), moving from LDA to GGA, to meta-GGA, and then to hybrid density functionals, PZ-SIC Perdew and Zunger (1981), etc., or by applying empirical corrections to the chosen functional, such as DFT+$U$. The HSE hybrid density functional Heyd et al. (2003) has become a standard for calculating band gaps in solids due to its reduced SIE; however, its computational cost—an order of magnitude higher than LDA, GGAs, and metaGGAs—limits its practicality for EPC calculations. DFT+$U$ offers an alternative to metaGGA functionals like r2SCAN, but the empirical nature of the $U$ parameters constrains its predictive power Anisimov et al. (1997).
Performance of r2SCAN in conventional superconductor MgB2Having examined the improvements that r2SCAN offers for predicting EPCs in complex compounds like CoO and NiO with $d$-electrons, we now demonstrate that r2SCAN also provides reliable predictions for EPC strength $\lambda$ [defined in Eq. (6)] in main group compounds. Here, we revisit the prototypical conventional superconductor MgB2, a well-known phonon-mediated multiband superconductor with a critical temperature ($T_{\text{c}}$) of approximately 39 K at ambient pressure Nagamatsu et al. (2001); Chen et al. (2001); Tsuda et al. (2001). MgB2 has been extensively studied, both experimentally Nagamatsu et al. (2001); Chen et al. (2001); Tsuda et al. (2001); Baron et al. (2004); d’Astuto et al. (2007a) and theoretically Choi et al. (2002a); Margine and Giustino (2013); Kortus et al. (2001); Yildirim et al. (2001); Choi et al. (2002b).
@@
Figure 3(c) presents the calculated phonon dispersion of MgB2 using r2SCAN and PBE functionals. Consistent with prior studies d’Astuto et al. (2007a); Yildirim et al. (2001), r2SCAN identifies four optical phonon modes at the $\Gamma$ point: $E_{1u}$ , $A_{2u}$, $E_{2g}$, and $B_{1g}$, with frequencies of 42.6, 50.2, 74.1, and 87.9 meV, respectively, in close agreement with experimental values d’Astuto et al. (2007a). Notably, the $E_{2g}$ mode, involving in-plane stretching of B atoms, is critical for EPCs in MgB2 Choi et al. (2002b); Margine and Giustino (2013); Choi et al. (2002a), as we can see in Figs. 3 (c) and (f). This comparison between r2SCAN and PBE results [Fig. 3(c)] shows excellent agreement for acoustic modes, while r2SCAN provides a higher and more accurate prediction for the $E_{2g}$ mode frequency of 74.1 meV, matching the experimental value of 75 meV obtained from Raman measurements d’Astuto et al. (2007a); Hlinka et al. (2001); Goncharov et al. (2001). On the other hand, PBE predicts a lower value of 63 meV for the $E_{2g}$ mode. This discrepancy with the experimental value is often attributed to the overestimated bond lengths by PBE, resulting in softer optical phonon modes. While previous studies have linked the difference between PBE predictions and experimental values to anharmonic effects Yildirim et al. (2001); Choi et al. (2002b), recent findings suggest that MgB2 exhibits only weak anharmonicity d’Astuto et al. (2007b). This suggests that the improved results by r2SCAN arise from a better description of exchange-correlation effects.
@@
Figure 3: Calculated electronic structure, phonon dispersion and EPC of MgB2. (a) The calculated band structure of MgB2 along high-symmetry lines in the BZ from r2SCAN (solid black lines) and PBE (dash gray lines). Orbital characters from r2SCAN are also shown, where Mg $s$, B $p_{x/y}$, and B $p_{z}$ orbitals are indicated in green, red, and blue, respectively. The splitting of $\pi$ and $\sigma$ bands from r2SCAN at the $M$ point is highlighted. Orbital projections from PBE are omitted for simplicity. (b) The total and partial density of states for each orbital using r2SCAN. (c) Phonon dispersion along high-symmetry lines in the BZ, comparing r2SCAN (solid black lines) and PBE (dashed gray lines). Marker color intensity corresponds to the magnitude of the EPC strength, $\lambda_{\textbf{q}\nu}$ from r2SCAN calculations. For simplicity, EPC strength $\lambda_{\textbf{q}\nu}$ values from PBE calculations are not shown. Experimental phonon data from inelastic x-ray scattering (magenta solid points), extracted from Ref. d’Astuto et al. (2007a), are also included for comparison. The measured frequency of the $E_{2g}$ mode at the $\Gamma$ point from Raman spectroscopy is 75 meV d’Astuto et al. (2007a); Hlinka et al. (2001); Goncharov et al. (2001), represented by a blue cross. (d) Calculated Eliashberg spectral functions $\alpha^{2}F(\omega)$ and cumulative EPC strength $\lambda$ for MgB2, using r2SCAN (red lines) and PBE (green lines). A k-mesh of $100\times 100\times 100$ and a q-mesh of $50\times 50\times 50$ were used for the BZ sampling. (e) Left panel: The calculated three-dimensional Fermi surface of MgB2 obtained using the Wannier model based on r2SCAN calculations with a $200\times 200\times 200$ k-mesh. Middle and right panels: The Fermi surface projections on the $k_{z}=0$ and $k_{z}=\pi/c$ planes, respectively. (f) Distribution of $k$-resolved EPC strength $\lambda_{\boldsymbol{k}}=\sum_{mn,\nu\boldsymbol{q}}|g_{mn,\nu}({\boldsymbol{k}},\boldsymbol{q})|^{2}\delta(\epsilon_{m\textbf{k}+\textbf{q}}-\epsilon_{\text{F}})\times(2/\hbar\omega_{\textbf{q}\nu})$ from r2SCAN across the $k_{x}-k_{y}$ plane at $k_{z}=0$ (left panel) and $k_{z}=\pi/c$ (right panel).
@@
Figure 3(d) shows the calculated isotropic Eliashberg function $\alpha^{2}F(\omega)$ of MgB2, based on Eq. (9), using the r2SCAN (red lines) and PBE (green lines) functionals. Our PBE results reproduce the dominant peak around 60 meV, primarily contributed by the E2g mode, consistent with previous findings Margine and Giustino (2013). A second peak appears around 91 meV, slightly higher than previous results obtained with the LDA functional Margine and Giustino (2013). The calculated $\alpha^{2}F(\omega)$ using the PBE functional yields an EPC strength $\lambda$ of approximately 0.70, close to the earlier value of 0.74 reported for the LDA functional Margine and Giustino (2013). The r2SCAN results, shown as a solid red line, exhibit similarities to the PBE results in the range of 0–50 meV, since the predicted phonon dispersions from both functionals are nearly identical for low-frequency phonons [Fig. 3 (c)]. However, the dominant peaks from the r2SCAN calculations appear in the range of 63–75 meV, which are much broader than those from the PBE results, reflecting the larger phonon bandwidth of the $E_{2g}$ mode predicted by r2SCAN. The second peak is also captured in our r2SCAN results. These results also reflect the softening of the optical phonon modes in PBE. The broader distribution of the dominant peak in $\alpha^{2}F(\omega)$ observed with r2SCAN results in an EPC strength $\lambda$ of 0.73, slightly larger than the PBE value of 0.70. According to Eqs. (10) and (11), the estimated $\omega_{log}$ is $\sim$58.1 meV, and the predicted $T_{\text{c}}$ ranges from 22.4$\sim$30.3 K, with the adjustable parameter $\mu^{*}$ set between 0.12 and 0.08, consistent with previous calculations Margine and Giustino (2013).
@@
Based on the above analysis, our results demonstrate that the r2SCAN functional provides a more reliable description of electronic structure, phonon dispersion, and EPC in MgB2 compared to PBE. While previous studies have shown that MgB2 is a multi-gap superconductor Tsuda et al. (2001); Chen et al. (2001); Choi et al. (2002a, b); Giustino et al. (2007); Margine and Giustino (2013); Floris et al. (2005), solving the superconducting gap equation is beyond the scope of the present work.
ConclusionOur study demonstrates the efficacy of the r2SCAN functional for accurately predicting EPC in exemplar transition-metal compounds CoO and NiO, as well as the conventional superconductor MgB2, along with improved descriptions of both electronic structures and phonon properties compared to standard DFT methods such as PBE. These improvements result primarily from the improved treatment of SIE in r2SCAN. With continuing improvements in the DFT functionals beyond SCAN and r2SCAN, we can expect parallel advances in the predictive power of ab initio EPC calculations and the associated physical properties across a wide range of materials, including systems with stronger electron-electron and electron-phonon interactions.
@@
MethodsFirst-principles calculations were performed using the projector-augmented-wave (PAW) method Kresse and Joubert (1999), as implemented in the Vienna ab initio simulation package (VASP) Kresse and Hafner (1993); Kresse and Furthmüller (1996). Both the r2SCAN and PBE density functionals were employed in our calculations. For calculating the EPC, we follow the Wannier-interpolation procedure outlined in Ref. Engel et al. (2020) while using the all-electron EPC matrix elements defined in Ref. Engel et al. (2022); Chaput et al. (2019). The Wannier representation of the electronic band structure was opained using Wannier90 Mostofi et al. (2014) with the help of the VASP2WANNIER90 interface. We calculate the EPC strength $\lambda$ and isotropic Eliashberg spectral function $\alpha^{2}F$, we have implemented the code based on the following equations.
@@
where $f_{n{\boldsymbol{k}}}$ is the Fermi occupation of the single-particle state $|\psi_{n\boldsymbol{k}}\rangle$ associated with the eigenvalue $\epsilon_{n\boldsymbol{k}}$, $w_{\boldsymbol{k}}$ are the weights of the dense ${\boldsymbol{k}}$-points, and $\eta$ indicates a real positive infinitesimal. We note here the spin index is omitted for simplicity. The EPC strength $\lambda$ is the full BZ average of the mode-resolved $\lambda_{\boldsymbol{q}\nu}$
@@
with $w_{\boldsymbol{q}}$ being the weights of the dense $\boldsymbol{q}$-points. In our implementations, the double delta approximation Marini et al. (2024) is adopted to calculate the imaginary part of the phonon self-energy, under which the Eq. 5 can be rewritten as follows:
@@
MgB2A high-energy cutoff of 520 eV was applied to truncate the plane-wave basis set. Crystal structures and ionic positions were fully optimized using a force convergence criterion of 0.001 eV/Å for each atom, along with a total energy tolerance of $10^{-8}$ eV. To relax the structures, we adopted a $\Gamma$-centered k-point mesh of $13\times 13\times 12$ to sample the primitive bulk BZ. For phonon calculations, a $6\times 6\times 6$ supercell with a $\Gamma$-centered k-point mesh of $7\times 7\times 9$ was employed using the finite displacement method. A Gaussian smearing with a small $\sigma$ of 0.01 was applied in our phonon calculations. Mg $p$ and $s$ states, along with B $p$ and $s$ states, were included in constructing the Wannier functions.
CoO and NiOAn initial $G$-type antiferromagnetic (AFM) structure with 4 atoms was used for structural relaxation. A high-energy cutoff of 600 eV was applied to truncate the plane-wave basis set. Crystal structures and ionic positions were fully optimized with a force convergence criterion of 0.001 eV/Å for each atom and a total energy tolerance of $10^{-8}$ eV. A $\Gamma$-centered k-point mesh of $15\times 15\times 15$ was employed for the structural relaxation. For phonon calculations, a $4\times 4\times 4$ supercell with a $\Gamma$-centered $k$-point mesh of $3\times 3\times 3$ was used, employing the finite displacement method. A Gaussian smearing with a small $\sigma$ of 0.01 was applied in our phonon calculations. Transition metal (Ni and Co) $d$ states and oxygen $p$ states were included in constructing the Wannier functions.

### v2

@@ Abstract
From the above analysis, accurate EPC predictions can be expected within the KS-DFT framework if the SIE in the exchange-correlation approximation is minimized. This can be achieved by advancing up Jacob’s ladder of DFT Perdew and Schmidt (2001), moving from LDA to GGA, to meta-GGA, and then to hybrid density functionals, PZ-SIC Perdew and Zunger (1981), etc., or by applying empirical corrections to the chosen functional, such as DFT+$U$ Anisimov et al. (1991, 1993). The HSE hybrid density functional Heyd et al. (2003) has become a standard for calculating band gaps in solids due to its reduced SIE; however, its computational cost—one or two orders of magnitude higher than LDA, GGAs, and metaGGAs—limits its practicality for EPC calculations. DFT+$U$ offers an alternative to metaGGA functionals like r2SCAN, but the empirical nature of the $U$ parameters reduces its predictive power Anisimov et al. (1991, 1993).
Performance of r2SCAN in conventional superconductor MgB2Having examined the improvements that r2SCAN offers for predicting EPCs in complex compounds like CoO and NiO with $d$-electrons, we now demonstrate that r2SCAN also provides reliable predictions of the EPC strength $\lambda$ [defined in Eq. (6)] in main group compounds. Here, we revisit the prototypical conventional superconductor MgB2, a well-known phonon-mediated multiband superconductor with a critical temperature ($T_{\text{c}}$) of approximately 39 K at ambient pressure Nagamatsu et al. (2001); Chen et al. (2001); Tsuda et al. (2001). MgB2 has been extensively studied, both experimentally Nagamatsu et al. (2001); Chen et al. (2001); Tsuda et al. (2001); Baron et al. (2004); d’Astuto et al. (2007a) and theoretically Choi et al. (2002a); Margine and Giustino (2013); Kortus et al. (2001); Yildirim et al. (2001); Choi et al. (2002b).
@@
Figure 3(c) presents the calculated phonon dispersion of MgB2 using the r2SCAN and PBE functionals. Consistent with prior studies d’Astuto et al. (2007a); Yildirim et al. (2001), r2SCAN identifies four optical phonon modes at the $\Gamma$ point: $E_{1u}$ , $A_{2u}$, $E_{2g}$, and $B_{1g}$, with frequencies of 42.6, 50.2, 74.1, and 87.9 meV, respectively, in close agreement with experimental values d’Astuto et al. (2007a). Notably, the $E_{2g}$ mode, involving in-plane stretching of B atoms, is critical for EPCs in MgB2 Choi et al. (2002b); Margine and Giustino (2013); Choi et al. (2002a), as we can see in Figs. 3 (c) and (f). This comparison between r2SCAN and PBE results [Fig. 3(c)] shows excellent agreement for acoustic modes, while r2SCAN provides a higher and more accurate prediction for the $E_{2g}$ mode frequency of 74.1 meV, matching the experimental value of 75 meV obtained from Raman measurements d’Astuto et al. (2007a); Hlinka et al. (2001); Goncharov et al. (2001). PBE predicts a lower value of 63 meV for the $E_{2g}$ mode. This discrepancy with the experimental value is often attributed to the overestimated bond lengths by PBE, resulting in softer optical phonon modes. While previous studies have linked the difference between PBE predictions and experimental values to anharmonic effects Yildirim et al. (2001); Choi et al. (2002b), recent findings suggest that MgB2 exhibits only weak anharmonicity d’Astuto et al. (2007b). This suggests that the improved results by r2SCAN arise from a better description of exchange-correlation effects.
@@
Figure 3: Calculated electronic structure, phonon dispersion and EPC of MgB2. (a) The calculated band structure of MgB2 along high-symmetry lines in the BZ from r2SCAN (solid black lines) and PBE (dash gray lines). Orbital characters from r2SCAN are also shown, where Mg $s$, B $p_{x/y}$, and B $p_{z}$ orbitals are indicated in green, red, and blue, respectively. The splitting of $\pi$ and $\sigma$ bands from r2SCAN at the $M$ point is highlighted. Orbital projections from PBE are omitted for clarity. (b) The total and partial density of states for each orbital using r2SCAN. (c) Phonon dispersion along high-symmetry lines in the BZ, comparing r2SCAN (solid black lines) and PBE (dashed gray lines). Marker color intensity corresponds to the magnitude of the EPC strength, $\lambda_{\textbf{q}\nu}$ from r2SCAN calculations. EPC strength $\lambda_{\textbf{q}\nu}$ values from PBE calculations are not shown. Experimental phonon data from inelastic X-ray scattering (magenta solid points), extracted from Ref. d’Astuto et al. (2007a), are also included for comparison. The measured frequency of the $E_{2g}$ mode at the $\Gamma$ point from Raman spectroscopy is 75 meV d’Astuto et al. (2007a); Hlinka et al. (2001); Goncharov et al. (2001), represented by a blue cross. (d) Calculated Eliashberg spectral functions $\alpha^{2}F(\omega)$ and cumulative EPC strength $\lambda$ for MgB2, using r2SCAN (red lines) and PBE (green lines). A k-mesh of $100\times 100\times 100$ and a q-mesh of $50\times 50\times 50$ were used for the BZ sampling. (e) Left panel: The calculated three-dimensional Fermi surface of MgB2 obtained using the Wannier model based on r2SCAN calculations with a $200\times 200\times 200$ k-mesh. Middle and right panels: The Fermi surface projections on the $k_{z}=0$ and $k_{z}=\pi/c$ planes, respectively. (f) Distribution of $k$-resolved EPC strength $\lambda_{\boldsymbol{k}}=\sum_{mn,\nu\boldsymbol{q}}|g_{mn,\nu}({\boldsymbol{k}},\boldsymbol{q})|^{2}\delta(\varepsilon_{m\textbf{k}+\textbf{q}}-\varepsilon_{\text{F}})\delta(\varepsilon_{n\textbf{k}}-\varepsilon_{\text{F}})\times(2/\omega_{\textbf{q}\nu})$ from r2SCAN across the $k_{x}-k_{y}$ plane at $k_{z}=0$ (left panel) and $k_{z}=\pi/c$ (right panel).
@@
Figure 3(d) shows the calculated isotropic Eliashberg function $\alpha^{2}F(\omega)$ of MgB2, based on Eq. (9), using the r2SCAN (red lines) and PBE (blue lines) functionals. Our PBE results reproduce the dominant peak around 60 meV, primarily contributed by the E2g mode, consistent with previous findings Margine and Giustino (2013). A second peak appears around 91 meV, slightly higher than previous results obtained with the LDA functional Margine and Giustino (2013). The calculated $\alpha^{2}F(\omega)$ using the PBE functional yields an EPC strength $\lambda$ of approximately 0.70, close to the earlier value of 0.74 reported for the LDA functional Margine and Giustino (2013). The r2SCAN results, shown as a solid red line, exhibit similarities to the PBE results in the range of 0–50 meV, since the predicted phonon dispersions from both functionals are nearly identical for low-frequency phonons [Fig. 3 (c)]. However, the dominant peaks from the r2SCAN calculations appear in the range of 63–75 meV, which are much broader than those from the PBE results, reflecting the larger phonon bandwidth of the $E_{2g}$ mode predicted by r2SCAN. The second peak at 91 meV is also captured in our r2SCAN results. These results also reflect the softening of the optical phonon modes in PBE. The broader distribution of the dominant peak in $\alpha^{2}F(\omega)$ observed with r2SCAN results in an EPC strength $\lambda$ of 0.73, slightly larger than the PBE value of 0.70. According to Eqs. (10) and (11), the estimated $\omega_{log}$ is $\sim$58.1 meV, and the predicted $T_{\text{c}}$ ranges from 22.4$\sim$30.3 K, with the adjustable parameter $\mu^{*}$ set between 0.12 and 0.08, consistent with previous calculations Margine and Giustino (2013).
@@
Based on the above analysis, our results demonstrate that the r2SCAN functional provides a reliable description of electronic structure, phonon dispersion, and EPC in MgB2 compared to PBE. While previous studies have shown that MgB2 is a multi-gap superconductor Tsuda et al. (2001); Chen et al. (2001); Choi et al. (2002a, b); Giustino et al. (2007); Margine and Giustino (2013); Floris et al. (2005), solving the superconducting gap equation is beyond the scope of the present work.
ConclusionCalculating EPCs for complex materials with $d$- and $f$-electrons has long been a challenge for first-principles methods due to the need for both accuracy and efficiency. In this study, we demonstrate that the efficient r2SCAN density functional significantly improves EPC predictions compared to the standard PBE density functional for transition-metal compounds such as CoO and NiO, alongside better descriptions of their electronic structures and phonon dispersions. Due to SIE, PBE incorrectly predicts CoO as metallic with imaginary phonon frequencies, rendering it unsuitable for EPC calculations. In contrast, r2SCAN reduces SIE, opens the band gap, stabilizes the phonon dispersion, and predicts strong EPCs for optical modes. Analysis of the dielectric screening tensors for CoO and NiO also indicates that these improvements stem from the reduction of SIE in r2SCAN compared to PBE.
Additionally, for the main-group compound MgB2, r2SCAN provides improved predictions of the band structure and phonon dispersion compared to PBE. Notably, r2SCAN predicts a phonon frequency for the critical E2g mode that closely matches Raman measurements, whereas PBE significantly underestimates this value. The resulting EPC strength $\lambda$ and isotropic Eliashberg spectral function $\alpha^{2}F(\omega)$ calculated using r2SCAN show good agreement with prior studies.
Our findings highlight the general efficacy of the r2SCAN functional in accurately predicting EPCs across a wide range of materials. The study underscores the potential of advanced functionals like r2SCAN to enhance the predictive power of first-principles EPC calculations, enabling deeper and more accurate insights into the physical properties of materials with strong electron-electron and electron-phonon interactions.
@@
MethodsFirst-principles calculations were performed using the projector-augmented-wave (PAW) method Kresse and Joubert (1999), as implemented in the Vienna ab initio simulation package (VASP) Kresse and Hafner (1993); Kresse and Furthmüller (1996). Both the r2SCAN and PBE density functionals were employed in our calculations. For calculating the EPC, we follow the Wannier-interpolation procedure outlined in Ref. Engel et al. (2020) while using the all-electron EPC matrix elements defined in Ref. Engel et al. (2022); Chaput et al. (2019). The Wannier representation of the electronic band structure was obtained using Wannier90 Mostofi et al. (2014) with the help of the VASP2WANNIER90 interface. We calculate the EPC strength $\lambda$ and the isotropic Eliashberg spectral function $\alpha^{2}F$ based on the following equations.
@@
where $f_{n{\boldsymbol{k}}}$ is the Fermi occupation of the single-particle state $|\psi_{n\boldsymbol{k}}\rangle$ associated with the eigenvalue $\epsilon_{n\boldsymbol{k}}$, $w_{\boldsymbol{k}}$ are the weights of the dense ${\boldsymbol{k}}$-points, and $\eta$ indicates a real positive infinitesimal. We note here that the spin index is omitted for simplicity. The EPC strength $\lambda$ is the full BZ average of the mode-resolved $\lambda_{\boldsymbol{q}\nu}$
@@
with $w_{\boldsymbol{q}}$ being the weights of the dense $\boldsymbol{q}$-points. In our implementations, the double delta approximation Marini et al. (2024) is adopted to calculate the imaginary part of the phonon self-energy, under which the Eq. (5) can be rewritten as follows:
@@
CoO and NiO All the calculations were performed within the density functional theory framework with projector-augmented wave potentials Kresse and Joubert (1999) using the VASP Kresse and Hafner (1993); Kresse and Furthmüller (1996). An initial $G$-type antiferromagnetic (AFM) structure with 4 atoms was used for structural relaxation. A high-energy cutoff of 600 eV was applied to truncate the plane-wave basis set. Crystal structures and ionic positions were fully optimized with a force convergence criterion of 0.001 eV/Å for each atom and a total energy tolerance of $10^{-8}$ eV. A $\Gamma$-centered k-point mesh of $15\times 15\times 15$ was employed for the structural relaxation. For phonon calculations, a $4\times 4\times 4$ supercell with a $\Gamma$-centered $k$-point mesh of $3\times 3\times 3$ was used, employing the finite displacement method. A Gaussian smearing with a small $\sigma$ of 0.01 was applied in our phonon calculations. Transition metal (Ni and Co) $d$ states and oxygen $p$ states were included in constructing the Wannier functions.
MgB2The calculations were performed within the density functional theory framework with projector-augmented wave potentials Kresse and Joubert (1999) using the VASP Kresse and Hafner (1993); Kresse and Furthmüller (1996). A high-energy cutoff of 520 eV was applied to truncate the plane-wave basis set. Crystal structures and ionic positions were fully optimized using a force convergence criterion of 0.001 eV/Å for each atom, along with a total energy tolerance of $10^{-8}$ eV. To relax the structures, we adopted a $\Gamma$-centered k-point mesh of $13\times 13\times 12$ to sample the primitive bulk BZ. For phonon calculations, a $6\times 6\times 6$ supercell with a $\Gamma$-centered k-point mesh of $7\times 7\times 9$ was employed using the finite displacement method. A Gaussian smearing with a small $\sigma$ of 0.01 was applied in our phonon calculations. Mg $p$ and $s$ states, along with B $p$ and $s$ states, were included in constructing the Wannier functions.